OpenStudy (anonymous):

The ratio of boys to girls is 3 to 5 in Art Club. In Music Club, there are 9 boys and has the same ratio of boys to girls in Art Club. How many girls are in Music Club? 18 girls 15 girls 5.4 girls

OpenStudy (anonymous):

15 girls

OpenStudy (anonymous):

thx :)

OpenStudy (anonymous):

3/5 9/x cross-multiply: 3x=45 45/3=15

OpenStudy (anonymous):

3/5 = 9/x 3x = 45 x = 45/3 x = 15 I agree